how do you evalute
-3/5+4

Answer by ewatrrr(24785)

 
Hi
Evaluating by combining Like terms:
-3/5 + 4 = -3/5 + 20/5 = 17/5 or 3 2/5


ewatrrr@aol.com
private/pay tutoring Inquiries only, please.